OpenStudy (carolyn.x):

An Olympic athlete of mass 80kg competes in a 100m race. What is The best estimate of his mean kinetic energy during the race? A 400J B 4000J C 40000J D 400000J can someone explain to me in getting the final answer?

OpenStudy (michele_laino):

hint: let's suppose that the time needed to the athlete to cover 100 meters is \(10\) seconds, then the average speed, is: \[\Large v = \frac{d}{t} = \frac{{100}}{{10}} = 10\quad m/\sec \] so the kinetic energy of the athlete, is: \[\Large KE = \frac{1}{2}m{v^2} = \frac{1}{2} \cdot 80 \cdot {10^2} = ...?joules\]